OpenStudy (anonymous):

The surface area of a square pyramid whose base side is 6 cm and height is 4 cm is _____ cm²

OpenStudy (austinl):



OpenStudy (austinl):



OpenStudy (austinl):

We have 4 triangles and a square. we need to find the side height of the triangles, do you know how we would do that?

OpenStudy (austinl):

Hello?

OpenStudy (anonymous):

I am so confused on what you already did.

OpenStudy (austinl):

I drew the pyramid, and labeled what we know. The vertical height is 4cm, and the side length is 6cm. So therefore, from the edge to the center is 3cm. Make sense?

OpenStudy (anonymous):

Kinda

OpenStudy (anonymous):

so slant hight is 5 base is 6 hight is 4 and edge to center is 3. so to find surface area I would brake it in to parts 4 triangles and a square?

OpenStudy (austinl):

Correct!

OpenStudy (austinl):

\[A_{sides}=2\times b\times h\] \[A_{bottom}=l\times w\]

OpenStudy (anonymous):

one triangle would be (1/2)(6)(5) and the square would be 6^2

OpenStudy (austinl):

Correct.

OpenStudy (anonymous):

and my answer would come out to 96

OpenStudy (austinl):

That is what I arrived at.

OpenStudy (anonymous):

ok thank you

OpenStudy (austinl):

You are very welcome! If you have any other questions I will be here for about another hour.